OpenStudy (anonymous):

write an equation for the line parallel to the line 6x=5 through point (9,-3)

OpenStudy (shamil98):

6x = 5, x = 5/6 this is just a coordinate, not a line.

OpenStudy (anonymous):

not necessary. it means its a straight line. you know those types

OpenStudy (anonymous):

would the answer just be x=5/6?

OpenStudy (mathmale):

No. x=5/6 is a vertical line with infinite/undefined slope. The new line has to be parallel to the old line and therefore is also a vertical line with infinite/undefined slope. This new line must pass thru the point (9,-3). Please write the equation of this new line, based upon this information.

OpenStudy (anonymous):

Would it be y=5/6x-3

OpenStudy (mathmale):

I explained in my previous post that this new line would have an infinite (or undefined) slope, so that contradicts your y=(5/6)x - 3, which has slope (5/6). So, my response to your question is "no."

OpenStudy (gorv):

@margejean parallel lines have same slope and same equation except their intercept or constant will be different 6x=5 0=6x-5 its parallel line will have equation 0=6x+c it is passing through (9,-3) so put this in equation and find c 0=6*9+c c=-54 so parallel line passine through (9,-3) will be 0=6x-54 6x=54

OpenStudy (gorv):

or solve it further 6x=54 x=9